# On the Number of Degenerate Simplex Pivots

### v2

@@ 1 Introduction
We stress here that the pivots considered in Theorem 1 are degenerate simplex pivots, meaning that each of them yields an improving direction at the given extreme point, though this direction is not feasible. It is important to point this out as in general, given two adjacent extreme points $x,x^{\prime}$, one can easily perform a sequence of basis exchanges that yield $x^{\prime}$ from $x$ by identifying the common tight linearly independent constraints, and introducing each of them to the current basis in any order, until the direction $x^{\prime}-x$ is seen. However, we here want a strategy that guarantees that each basis exchange will define an improving direction.
We then discuss some byproducts of our result in Section 3. In particular, we revise the analysis of the simplex algorithm by Kitahara and Mizuno [19, 20] who bound the number of non-degenerate pivots it in terms of $n,m$ and the maximum and the minimum non-zero coordinate of a basic feasible solution. Their analysis combined with our degeneracy-escaping technique show that the total number of simplex pivots (both degenerate and non-degenerate) can be bounded in a similar way. As a consequence, one can have a strongly-polynomial number of simplex pivots for several combinatorial LPs (see observations following Corollary 6).
@@ 2 Antistalling pivot rule
The goal of this section is to prove Theorem 1. Before doing that, we state some preliminaries, and also give a geometric intuition behind the result.
@@ 2.1 Preliminaries
For $A\in\mathrm{\hskip 0.28453ptIR}^{m\times n}$ and $B\subseteq[n]$ we use $A_{B}$ to denote the submatrix of $A$ formed by the columns indexed by $B$. Similarly, for $x\in\mathrm{\hskip 0.28453ptIR}^{n}$ and $B\in[n]$, the notation $x_{B}$ is used to denote a vector consisting of the entries of $x$ indexed by $B$. A basis of (1) is a subset $B\subseteq[n]$ with $|B|=m$ and $A_{B}$ being non-singular. The point $x$ with $x_{B}=A_{B}^{-1}b$, $x_{N}=\mathbbm{O}$ where $N:=[n]\setminus B$ is a basic solution of (1) with basis $B$. If additionally $x_{B}\geq 0$, both the basic solution $x$ and the basis $B$ are feasible. If $x_{i}>0$ for all $i\in B$, then $B$ and $x$ are called non-degenerate and degenerate otherwise, i.e., if $x_{i}=0$ for some $i\in B$. We let $\overline{A}:=A_{B}^{-1}A_{N}$ and $\overline{c}_{N}^{T}:=c_{N}^{T}-c_{B}^{T}\overline{A}$. In particular, $\overline{c}_{N}\in\mathrm{\hskip 0.28453ptIR}^{N}$ is the vector of so-called reduced costs for the basis $B$. The coordinates of the reduced cost vector will be addressed as $\overline{c}_{N,i}$, where the first subscript will be dropped if the basis $B=[n]\setminus N$ is clear from the context.
@@ 2.2 A geometric intuition
2.3 Proof of Theorem 1
Proof.
@@ 1 $r$ $k$ $m$ $f$ $m+t$ $n$
(*$(*)\quad\min_{i\in Eq^{>}(f)}\frac{x_{i}}{\overline{A}_{if}}>0\mbox{ where }Eq^{>}(f):=\{i\in[m]\mid\overline{A}_{if}>0\}\,.$
In particular, this means $Eq^{>}(f)\subseteq[m]\setminus[k]$. We perform a non-degenerate pivot, by selecting an index that minimizes (*)(*)( * ) as the leaving index.
Case II: $\overline{A}_{if}>0$ for some $i\in[k]\setminus[r]$. In this case, we perform a degenerate pivot by selecting $i$ as the leaving index. Let $B^{\prime}:=B\cup\{f\}\setminus\{i\}$. Because of degeneracy, the basic solution associated with $B^{\prime}$ is still $x$, and hence $y^{0}$ is still an improving feasible direction for $x$. Note that $|Q_{2}(y^{0},B^{\prime})|=|Q_{2}(y^{0},B)|-1$— since $y^{0}_{f}>0$ but $y^{0}_{i}=0$ by definition. Repeat the same for $B^{\prime}$ and $y^{0}$.
@@
The key observation is that when the third case of the above case distinction has occurred, repeating the same for $B$ and $y^{1}$ falls into the second case (because the basis $B$ has not changed, meanwhile $y^{1}_{g}=x_{g}=0$ with $g\in B$ and $\overline{A}_{gf}>0$). Therefore, after each pivot, the cardinality of the support of the improving feasible direction in the non-basic indices decreases. Hence, a sequence of $|Q_{2}(y,B)|$ degenerate pivots would yield an improving feasible direction $y^{\prime}$ and a basis $B^{\prime}$ of $x$ with $Q_{2}(y^{\prime},B^{\prime})=0$, which in turn implies $0=\overline{c}^{T}_{N^{\prime}}y^{\prime}_{N^{\prime}}=c^{T}y^{\prime}$ due to (2), yielding a contradiction. Hence the number of consecutive degenerate pivots with the suggested pivot rule cannot exceed $n-m-1$.
@@ 3 Exploiting the antistalling rule for general bounds
Now, combining Lemma 2 and Lemma 4 allows to conclude that the simplex algorithm with the antistalling pivot rule described in Section 2 encounters at most $(n-m)\lceil\frac{(n-m)\Delta}{\delta}\log_{e}\big{(}m\frac{\Delta}{\delta}\big{)}\rceil$ distinct basic feasible solutions. Since Theorem 1 in turn bounds the number of consecutive non-degenerate steps, combining it with the latter result yields a bound on the total number of pivots required to reach an optimal vertex. However, this bound does not take into account the number of (degenerate) pivots that might have to be performed at an optimal vertex to encounter a basis satisfying the optimality criterion. This is due to the fact that the antistalling pivot rule requires an improving feasible direction, which we do not have at an optimal vertex. Hence, we have to handle this case separately.
@@
Let $B$ and $B^{\star}$ be a non-optimal and an optimal basis, both associated with an optimal basic feasible solution $x^{\star}$ of (1). Since $B$ is not optimal, there exists $f\in N$ with $\overline{c}_{N,f}=c^{T}z<0$ with $z$ as in (3). Observe that there exists $i\in B\cap N^{\star}$ with $\overline{A}_{if}>0$ and $x_{i}=0$: otherwise, all coordinates of $z_{N^{\star}}$ would be non-negative and hence $z\in C_{N^{\star}}:=\{x\in\mathrm{\hskip 0.28453ptIR}^{n}\mid Ax=0,x_{N^{\star}}\geq 0\}$. The latter however contradict the fact that all extreme rays of the above cone $C_{N^{\star}}$ have non-negative scalar products $\overline{c}_{N^{\star}}$ with $c$ due to optimality of $B^{\star}$. Hence one could perform a simplex pivot on $B$ with entering variable $f$ and leaving variable $i$. Let $B^{\prime}:=\big{(}B\setminus\{i\}\big{)}\cup\{f\}$ and $N^{\prime}:=\big{(}N\setminus\{f\}\big{)}\cup\{i\}$. Note that $\overline{c}_{N^{\prime},i}\geq 0$ and $i\in N^{\prime}\cap N^{\star}$. Either $B^{\prime}$ is a optimal basis and we stop, or there exists $j\in N^{\prime}\setminus\{i\}$ with $\overline{c}_{N^{\prime},j}<0$. In the latter case, however, we can enforce the constraint $x_{i}=0$ by removing the variable $x_{i}$ together with the corresponding column of $A$ and entry $c_{i}$ of $c$ from (1). By doing so we obtain a new LP with the number of variables smaller by one that has $B^{\prime}$ and $B^{\star}$ as a non-optimal and an optimal basis, respectively, since $\overline{c}_{N^{\prime}\setminus\{i\},j}<0$ and $\overline{c}_{N^{\star}\setminus\{i\}}\geq 0$. We can set $B=B^{\prime}$ and repeat the process. Since at each iteration a variable with index from $N^{\star}$ gets removed, after at most $n-m$ iterations $B=B^{\star}$. Since the number of degenerate pivots at an optimal vertex is bounded by $n-m$, we can state the following result.
Theorem 5.
@@
Given an LP (1) and an initial feasible basis, there exists a pivot rule that makes the simplex algorithm reach an optimal basis in at most $(n-m)^{2}\lceil\frac{(n-m)\Delta}{\delta}\log_{e}\big{(}m\frac{\Delta}{\delta}\big{)}\rceil$ simplex pivots.
@@
In general, the value of $\delta$ is NP-hard to compute [22]. However, observe that for integral $A$ and $b$ (which can be assumed without loss of generality for rational LPs), one can bound $\Delta\leq||b||_{1}\Delta_{A}$ and $\delta\geq\frac{1}{\Delta_{A}}$ due to Cramer’s rule, where $\Delta_{A}$ is the largest absolute value of the determinant of a square $m\times m$ submatrix of $A$. Then the next statement a straightforward corollary of the above theorem.
Corollary 6.
@@
For any basic feasible solution of an LP (1) with integral $A$ and $b$, there exists a sequence of at most $(n-m)^{2}\lceil(n-m)\Delta_{A}^{2}||b||_{1}\log_{e}\big{(}m\Delta_{A}^{2}||b||_{1}\big{)}\rceil$ simplex pivots leading to an optimal basis.
@@
We conclude this section by observing that, using the above corollary, one can prove existence of short sequences of simplex pivots (that is, of length strongly-polynomial in $n,m$) between any two extreme points of several combinatorial LPs. Just to mention a few:
@@
(a) LPs modeling matching/vertex-cover/edge-cover/stable-set problems in bipartite graphs. For matchings, the LP maximizes a linear function over a set of constraint of the form $\{A^{\prime}x\leq 1,x\geq 0\}$, where the coefficient matrix $A^{\prime}$ is the node-edge incidence matrix of an undirected bipartite graph. After putting the LP in standard equality form by adding slack variables, we get constraints of the form $\{Ax=1,x\geq 0\}$, where $A$ is a totally unimodular matrix (and so $\Delta_{A}=1$). The result then follows from Corollary 6. The same holds for vertex-cover (constraints of the form $\{{A^{\prime}}^{T}x\geq 1,x\geq 0\}$), edge-cover (constraints of the form $\{{A^{\prime}}x\geq 1,x\geq 0\}$), stable-set (constraints of the form $\{{A^{\prime}}^{T}x\leq 1,x\geq 0\}$).
@@
(b) LP relaxations for the above problems for general graphs (i.e., optimization over the fractional matching/vertex-cover/edge-cover/stable-set polytopes). The same LPs as in (a), in general graphs, have a constraint matrix $A^{\prime}$ (resp. ${A^{\prime}}^{T}$) that is not totally unimodular. However, the set of constraints defines half-integral polytopes (see [2, 27]). Note that, after putting the LPs in standard equality form, the slack variables can be loosely bounded from above by the number of variables $n$. Hence, $\Delta\leq n$ and $\delta\geq\frac{1}{2}$, and the result follows from Theorem 5.
@@
(c) LP for the stable marriage problem [30, 38]. The (exact) LP formulation has constraints of the form $\{A^{\prime}x\leq 1,B^{\prime}x\geq 1,x\geq 0\}$. After putting the LP in standard equality form, the slack variables can be bounded by $1$. Hence, $\Delta=\delta=1$, and the result follows from Theorem 5.
@@
(d) LPs modeling various flow problems (such as max flows, min cost flows, flow circulations) with unit (or bounded) capacity values. For the corresponding LPs (see [32]), the constraint matrix will be a node-arc incidence matrix of a directed graph, which is totally unimodular, and the right-hand-side vector will be bounded by the capacity values. Similarly to (a), one can rely on Corollary 6 to get the result.

### v3

@@ Abstract
Keywords:
Simplex Algorithm Linear Programming Degeneracy.
@@ 1 Introduction
We stress here that the pivots considered in Theorem 1.1 are degenerate simplex pivots, meaning that each of them yields an improving direction at the given extreme point, though this direction is not feasible. It is important to point this out as in general, given two adjacent extreme points $x,x^{\prime}$, one can easily perform a sequence of basis exchanges that yield $x^{\prime}$ from $x$ by identifying the common tight linearly independent constraints, and introducing each of them to the current basis in any order, until the direction $x^{\prime}-x$ is seen. However, we here want a strategy that guarantees that each basis exchange is realizable by the simplex algorithm and hence defines an improving direction.
We then discuss some byproducts of our result in Section 3. In particular, we revise the analysis of the simplex algorithm by Kitahara and Mizuno [21, 22] who bound the number of non-degenerate pivots in terms of $n,m$ and the maximum and the minimum non-zero coordinate of a basic feasible solution. Their analysis combined with our degeneracy-escaping technique show that the total number of simplex pivots (both degenerate and non-degenerate) can be bounded in a similar way. As a consequence, one can have a strongly-polynomial number of simplex pivots for several combinatorial LPs. In addition, we perform some computational experiments to evaluate the performance of the antistalling pivot rule in practice, reported in Section 4.
@@ 2 Antistalling pivot rule
The goal of this section is to prove Theorem 1.1. Before doing that, we state some preliminaries, and also give a geometric intuition behind the result.
@@ 2.1 Preliminaries
For $A\in\mathrm{\hskip 0.28453ptIR}^{m\times n}$ and $B\subseteq[n]$ we use $A_{B}$ to denote the submatrix of $A$ formed by the columns indexed by $B$. Similarly, for $x\in\mathrm{\hskip 0.28453ptIR}^{n}$ and $B\in[n]$, the notation $x_{B}$ is used to denote a vector consisting of the entries of $x$ indexed by $B$. A basis of (1) is a subset $B\subseteq[n]$ with $|B|=m$ and $A_{B}$ being non-singular. The point $x$ with $x_{B}=A_{B}^{-1}b$, $x_{N}=\mathbb{0}$ where $N:=[n]\setminus B$ is a basic solution of (1) with basis $B$. If additionally $x_{B}\geq 0$, both the basic solution $x$ and the basis $B$ are feasible. If $x_{i}>0$ for all $i\in B$, then $B$ and $x$ are called non-degenerate and degenerate otherwise, i.e., if $x_{i}=0$ for some $i\in B$. We let $\overline{A}:=A_{B}^{-1}A_{N}$ and $\overline{c}_{N}^{T}:=c_{N}^{T}-c_{B}^{T}\overline{A}$. In particular, $\overline{c}_{N}\in\mathrm{\hskip 0.28453ptIR}^{N}$ is the vector of so-called reduced costs for the basis $B$. The coordinates of the reduced cost vector will be addressed as $\overline{c}_{N,i}$, where the first subscript will be dropped if the basis $B=[n]\setminus N$ is clear from the context.
@@ 2.2 A geometric intuition
2.3 Proof of Theorem 1.1
Proof
@@ 1 $r$ $k$ $m$ $f$ $m+t$ $n$
$(*)\quad\min_{i\in Eq^{>}(f)}\frac{x_{i}}{\overline{A}_{if}}>0\mbox{ where }Eq^{>}(f):=\{i\in[m]\mid\overline{A}_{if}>0\}\,.$
In particular, this means $Eq^{>}(f)\subseteq[m]\setminus[k]$. We perform a non-degenerate pivot, by selecting an index that minimizes ()(*)( ∗ ) as the leaving index.
Case II: $\overline{A}_{if}>0$ for some $i\in[k]\setminus[r]$. In this case, we perform a degenerate pivot by selecting $i$ as the leaving index. Let $B^{\prime}:=B\cup\{f\}\setminus\{i\}$. Because of degeneracy, the basic solution associated with $B^{\prime}$ is still $x$, and hence $y^{0}$ is still an improving feasible direction for $x$. Note that $|Q_{2}(y^{0},B^{\prime})|=|Q_{2}(y^{0},B)|-1$ since $y^{0}_{f}>0$ but $y^{0}_{i}=0$ by definition. Repeat the same for $B^{\prime}$ and $y^{0}$.
@@
The key observation is that when the third case of the above case distinction has occurred, repeating the same for $B$ and $y^{1}$ falls into the second case (because the basis $B$ has not changed, meanwhile $y^{1}_{g}=x_{g}=0$ with $g\in B$ and $\overline{A}_{gf}>0$). Therefore, after each degenerate pivot, the cardinality of the support of the improving feasible direction in the non-basic indices decreases. Hence, a sequence of $|Q_{2}(y^{0},B)|$ degenerate pivots would yield an improving feasible direction $y^{\prime}$ and a basis $B^{\prime}$ of $x$ with $Q_{2}(y^{\prime},B^{\prime})=0$, which in turn implies $0=\overline{c}^{T}_{N^{\prime}}y^{\prime}_{N^{\prime}}=c^{T}y^{\prime}$ due to (2), yielding a contradiction. Hence the number of consecutive degenerate pivots with the suggested pivot rule cannot exceed $n-m-1$. ∎
We conclude with a remark which will be useful in the next section.
Remark 1
Let $x$ be the currently considered basic feasible solution with basis $B$. If $y^{0}$ is chosen to be equal to $\tilde{x}-x$ for some basic feasible solution $\tilde{x}$ of our LP with $c^{T}\tilde{x}>c^{T}x$, then one can observe that $|Q_{2}(y^{0},B)|\leq m$ because $\tilde{x}$ has at most $m$ non-zero coordinates. By the proof of Theorem 1.1 above, the latter value strictly decreases with each degenerate step produced by the antistalling pivot rule. Hence, the total number of consecutive degenerate pivots at the vertex $x$ can be strengthen to be at most $\min\{n-m-1,m-1\}$.
@@ 3 Exploiting the antistalling rule for general bounds
Now, combining Lemma 1 and Lemma 3 allows to conclude that the simplex algorithm with the antistalling pivot rule described in Section 2 encounters at most $(n-m)\lceil\frac{(n-m)\Delta}{\delta}\log_{e}\big{(}m\frac{\Delta}{\delta}\big{)}\rceil$ distinct basic feasible solutions. Since Theorem 1.1 and Remark 1 in turn bound the number of consecutive non-degenerate steps by $\min\{n-m-1,m-1\}$ (again, due to the choice of the improving feasible direction), combining it with the latter result yields a bound on the total number of pivots required to reach an optimal vertex. However, this bound does not take into account the number of (degenerate) pivots that might have to be performed at an optimal vertex to encounter a basis satisfying the optimality criterion. This is due to the fact that the antistalling pivot rule requires an improving feasible direction, which we do not have at an optimal vertex. Hence, we have to handle this case separately.
@@
Let $B$ and $B^{\star}$ be a non-optimal and an optimal basis, both associated with an optimal basic feasible solution $x^{\star}$ of (1). Assume $B$ is the basis of $x^{\star}$ obtained by the antistalling pivot rule. Since $B$ is not optimal, there exists $f\in N$ with $\overline{c}_{N,f}=c^{T}z<0$ with $z$ as in (3). Observe that there exists $i\in B\cap N^{\star}$ with $\overline{A}_{if}>0$ and $x_{i}=0$: otherwise, all coordinates of $z_{N^{\star}}$ would be non-negative and hence $z\in C_{N^{\star}}:=\{x\in\mathrm{\hskip 0.28453ptIR}^{n}\mid Ax=0,x_{N^{\star}}\geq 0\}$. The latter however contradict the fact that all extreme rays of the above cone $C_{N^{\star}}$ have non-negative scalar products $\overline{c}_{N^{\star}}$ with $c$ due to optimality of $B^{\star}$. Hence one could perform a simplex pivot on $B$ with entering variable $f$ and leaving variable $i$. Let $B^{\prime}:=\big{(}B\setminus\{i\}\big{)}\cup\{f\}$ and $N^{\prime}:=\big{(}N\setminus\{f\}\big{)}\cup\{i\}$. Note that $\overline{c}_{N^{\prime},i}\geq 0$ and $i\in N^{\prime}\cap N^{\star}$. Either $B^{\prime}$ is a optimal basis and we stop, or there exists $j\in N^{\prime}\setminus\{i\}$ with $\overline{c}_{N^{\prime},j}<0$. In the latter case, however, we can enforce the constraint $x_{i}=0$ by removing the variable $x_{i}$ together with the corresponding column of $A$ and entry $c_{i}$ of $c$ from (1). By doing so we obtain a new LP with the number of variables smaller by one that has $B^{\prime}$ and $B^{\star}$ as a non-optimal and an optimal basis, respectively, since $\overline{c}_{N^{\prime}\setminus\{i\},j}<0$ and $\overline{c}_{N^{\star}\setminus\{i\}}\geq 0$. We can set $B=B^{\prime}$ and repeat the process. Since at each iteration a variable with index from $N^{\star}$ gets removed, after at most $n-m$ iterations $B=B^{\star}$. Since the number of degenerate pivots at an optimal vertex is bounded by $n-m$, we can state the following result.
Theorem 3.1
@@
Given an LP (1) and an initial feasible basis, there exists a pivot rule that makes the simplex algorithm reach an optimal basis in at most
$\min\{n-m,m\}(n-m)\lceil\frac{(n-m)\Delta}{\delta}\log_{e}\big{(}m\frac{\Delta}{\delta}\big{)}\rceil$
simplex pivots.
@@
In general, the value of $\delta$ is NP-hard to compute [24]. However, observe that for integral $A$ and $b$ (which can be assumed without loss of generality for rational LPs), one can bound $\Delta\leq||b||_{1}\Delta_{A}$ and $\delta\geq\frac{1}{\Delta_{A}}$ due to Cramer’s rule, where $\Delta_{A}$ is the largest absolute value of the determinant of a square submatrix of $A$. Then the next statement a straightforward corollary of the above theorem.
Corollary 1
@@
For any basic feasible solution of an LP (1) with integral $A$ and $b$, there exists a sequence of at most
$\min\{n-m,m\}(n-m)\lceil(n-m)\Delta_{A}^{2}||b||_{1}\log_{e}\big{(}m\Delta_{A}^{2}||b||_{1}\big{)}\rceil$
simplex pivots leading to an optimal basis.
3.1 Application to combinatorial LPs
@@
We conclude this section by observing that, using the latter corollary, one can prove the existence of short sequences of simplex pivots (that is, of length strongly-polynomial in $n,m$) between any two extreme points of several combinatorial LPs, that is, LPs modeling the set of feasible solutions of famous combinatorial optimization problems. We report a few examples below.
(a) LPs modeling matching/vertex-cover/edge-cover/stable-set problems in bipartite graphs.
@@
For matchings, the LP maximizes a linear function over a set of constraint of the form $\{A^{\prime}x\leq 1,x\geq 0\}$, where the coefficient matrix $A^{\prime}$ is the node-edge incidence matrix of an undirected bipartite graph. After putting the LP in standard equality form by adding slack variables, we get constraints of the form $\{Ax=1,x\geq 0\}$, where $A$ is a totally unimodular matrix (and so $\Delta_{A}=1$). The result then follows from Corollary 1. The same holds for vertex-cover (minimizing a linear function over constraints of the form $\{{A^{\prime}}^{T}x\geq 1,x\geq 0\}$), edge-cover (minimizing a linear function over constraints of the form $\{{A^{\prime}}x\geq 1,x\geq 0\}$), stable-set (maximizing a linear function over constraints of the form $\{{A^{\prime}}^{T}x\leq 1,x\geq 0\}$).
(b) LPs modeling optimization over the fractional matching/vertex-cover/edge-cover/stable-set polytopes.
@@
These correspond to the natural LP relaxations of the problems discussed in (a), for general graphs. The same LPs as in (a), in non-bipartite graphs, have a constraint matrix $A^{\prime}$ (resp. ${A^{\prime}}^{T}$) that is not totally unimodular. However, the set of constraints defines half-integral polytopes (see [2, 29]). Note that, after putting the LPs in standard equality form, the slack variables can be loosely bounded from above by the number of variables $n$. Hence, $\Delta\leq n$ and $\delta\geq\frac{1}{2}$, and the result follows from Theorem 3.1.
(c) LP for the stable marriage problem.
The classical stable marriage problem is defined on a bipartite graph where each node has a strict preference order over the set of its neighbours. One looks for a matching that does not contain any blocking pair, that is, a pair of nodes that mutually prefer each other with respect to their matched neighbour. There is an (exact) LP formulation for the problem [33, 42], that has constraints of the form $\{A^{\prime}x\leq 1,B^{\prime}x\geq 1,x\geq 0\}$. Here $A^{\prime}$ is again the node-edge incidence matrix of an undirected bipartite graph, while $B^{\prime}$ is a matrix that stems from imposing an additional constraint for each edge $\{uv\}$, that essentially prevents $\{uv\}$ from being a blocking pair:
$x_{uv}+\sum_{w:w>_{u}v}x_{uw}+\sum_{w:w>_{v}u}x_{vw}\geq 1$
@@
In the above expression, $w>_{u}v$ (resp. $w>_{v}u$) means that $u$ prefers $w$ over $v$ (resp. $v$ prefers $w$ over $u$). After putting the LP in standard equality form, the slack variables can be bounded by $1$. Hence, $\Delta=\delta=1$, and the result follows from Theorem 3.1.
@@
(d) LPs modeling various flow problems (such as max flows, min cost flows, flow circulations) with unit (or bounded) capacity values.
Flow problems in capacitated graphs are modeled using LPs with constraints of the form $\{{A^{\prime}}x=b,l\leq x\leq u\}$ where the constraint matrix $A^{\prime}$ is here a node-arc incidence matrix of a directed graph, which is totally unimodular (see [35]). The right-hand-side vectors ($b,l,u$) can be bounded in terms of the total capacity values. Therefore, assuming these are bounded integers, one can rely on Corollary 1 to get the result, similarly to (a).
One can compute the corresponding pivot sequence for the problems mentioned in (a)-(d) by running the simplex with the antistalling pivot rule. However, one needs to presolve the LP and obtain an optimal basic solution in order to use it to guide the antistalling pivot rule.
We remark that the existence of a strongly polynomial sequence of pivots for general min-cost flows (hence for problems in (d)), was known already, as it follows from the fundamental work of Orlin [31], that applies even in the case of large capacity values. In comparison, our bound is weaker as it requires the capacity values to be bounded.
Instead, despite being known to be solvable in strongly polynomial time, to the best of our knowledge the problems in (a)-(c) were not previously known to admit strongly polynomial bounds on the number of simplex pivots for their natural LP formulations.
4 Computational experiments
In this section we report some computational experiments that we conducted to evaluate the performance of the antistalling pivot rule when applied to actual linear programs.
For the implementation, we used the python package CyLP [41] which wraps COIN-OR’s CLP solver [9] and provides tools for implementing a preferred pivot rule in python to substitute CLP’s built-in ones. For the tests, we used the benchmark Netlib LP dataset [30] containing 93 LPs of diverse dimensions and sparsity. All our experiments were conducted on a laptop with Intel Core i7-13700H 2.40GHz CPU and 32GB RAM.
For each of the aforementioned LPs, we ran the simplex algorithm with our antistalling pivot rule, and in addition with other 5 well-known pivot rules (which [41] provides implementations for). These are:
  • Dantzig’s rule: the entering variable is the one with the most negative reduced cost;
  • Steepest edge: the entering variable is the one which yields a direction $z$ maximizing $\frac{-c^{T}z}{||z||_{1}}$ (i.e., maximizing the improvement normalized according to the 1-norm);
  • LIFO: the entering variable is the one that minimizes the number of iterations that have past since the variable left the basis;
  • Most Frequent: the entering variable is the one that maximizes the number of times it was previously chosen as the entering variable;
  • Bland’s rule: the entering variable is the one with the lowest index.
We tracked the number of pivots required by each of the above rules. For our antistalling pivot rule, we guided it using the feasible direction $y=x^{\star}-x$ at any non-optimal vertex $x$ of the LP, for a pre-computed optimal basic solution $x^{*}$ (in particular, the one founded by Dantzig’s pivot rule). Intuitively, this represents the best direction that can possibly guide our rule, as it is the direction leading immediately to an optimal solution. Our objective with the computational experiments is to see whether this choice translates into few pivots in practice, since $y$ might change during the degenerate steps and therefore in reality we do not have control on the actual edge-direction we end up moving along.
Among the 93 instances in the dataset, there were 84 LPs for which each of the pivot rules was able to find an optimal solution with 30 minutes of timeout. We restrict our report to these LPs. For the sake of graphical representation, we further divided these 84 problems into 2 groups of 42 LPs each, by considering the maximum number of iterations required by any of the 6 tested pivot rules. In particular, the first group contains 42 LPs for which all pivot rules were able to arrive to an optimal solution within 2369 pivots. The second group contains the remaining 42 LPs. Figure 2 shows the results.
Figure 2: Both the upper and the lower plots show the number of iterations required to solve the LPs by each of the 6 considered pivot rules. The 42 LPs in the upper plot are the ones solved within 2369 pivots by all pivot rules. The remaining 42 LPs are considered in the lower plot. Each of the so-called violin plot visualises the distribution of the corresponding 42 numbers. White dots depict means, thick black dashes represent the intervals between 0.25- and 0.75-quantiles.
The computational experiments show that the antistalling pivot rule, guided by a known optimal solution, actually performs quite well, and most often manages to find a relatively short sequence of pivots to an optimal basis compared to other pivot rules. We highlight in particular the problem 25fv47, where any other pivot rule required at least 7892 pivots whereas the antistalling only needed 1468. On the other hand, there are instances where the antistalling pivot rule actually showed the worse performance compared to all its competitors. Those are the LPs fit2p and seba: 69711 and 976 pivots against at most 61925 and 873, respectively.
5 Final Remarks
We would like to conclude the paper with some remarks and potential research directions that originate from our findings.
The results of this work show the existence of a short sequence of degenerate simplex pivots that leads out of degeneracy. It would be interesting to show that similar results can be achieved through perturbations. Informally, perturbing the right-hand-side of the constraints has the effect of “splitting” a degenerate vertex into a set of (possibly exponentially many) distinct vertices, each associated to a distinct basis. It is legitimate to ask whether every new vertex allows for a short monotone path (where monotonicity is with respect to the linear objective function of the LP), which leads to a vertex out of this set. Note that the existence of the short degeneracy-escaping sequence of simplex pivots showed in this work, generally does not answer the latter question since some bases used in the sequence might become infeasible after perturbation.
Next, we would like to connect the above question to the recent work [20]. Given a basis $B$ defining a degenerate vertex $v$, one could imagine to separate $v$ from all its neighbouring vertices by a hyperplane: this way, one obtain a pyramid (with $v$ at its apex). Suppose this pyramid is full-dimensional and is described by some $Ax\leq b$: then, the above question is equivalent to perturbing the side facets of the pyramid to allow for a short path from the new apex defined by the given basis $B$, to any of the vertices in the bottom of the pyramid. This can be done relying on the techniques used for the construction of a so-called rock extension [20] (when viewing the base of the pyramid as the $d-1$-dimensional original polytope, and the pyramid as its $d$-dimensional extension). Such rock extension has been proven to have the property that the apex is connected to any bottom vertex by a path of at most linear length. However these paths are not necessarily monotone for a given linear objective function, and moreover, the construction of the rock extension requires not only perturbing the right-hand-side $b$ but the left-hand-side $A$ as well.
Another future research direction, that seems to be supported also by the performed computational experiments, would be to analyze the relation between the improving feasible direction $y$ that is required by our antistalling pivot rule, and the actual edge-direction $z$ along which one ends up moving after all the degenerate pivots. It would be very interesting to identify some conditions which ensure that $z$ is a good approximation of $y$, e.g., in terms of objective function’s improvement.
Futhermore, as it was mentioned in Section 2, in order for our bound to hold, the only requirement we need to impose when performing a pivot is that the entering variable lies in the support of $y$ (besides of course having negative reduced cost). In order to obtain the results from Section 3, where both degenerate and non-degenerate pivots have to be taken into account, we chose the entering variable to be the one with the most negative reduced cost, i.e., we used Dantzig’s rule just restricted to the coordinates in $\mathrm{supp}(y)$. Of course, one could think of analyzing the performance of the antistalling pivot rule with other choices of entering coordinates from $\mathrm{supp}(y)$, e.g., according to steepest edge or shadow vertex. Variations of these latter two rules, in particular, have been shown to play a key role in the context of 0/1 polytopes [4].